OpenStudy (anonymous):

[9.05] What are the exact solutions of x2 − 3x − 7 = 0? Medal to best answer

OpenStudy (anonymous):

@JackJones

OpenStudy (johnweldon1993):

Have you heard of the quadratic formula?

OpenStudy (anonymous):

No I haven't

OpenStudy (johnweldon1993):

Okay...what you have....is an equation x² - 3x -7 = 0 With 3 coefficients 1x² (1 is the coefficient) -3x (-3 is the coefficient) and -7 (7 is the coefficient) assign these to letter 1 = a -3 = b -7 = c Use this equation: \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] *hope that shows up...didn't load correctly* but plug each of those numbers in for the letters they correspond to in that equation.....This will give you 2 answers....understandable?

OpenStudy (anonymous):

Thank you for helping me understand!!

OpenStudy (johnweldon1993):

no problem man!